OpenStudy (anonymous):

Solve the equation x^2 − 36 = 0. Select one of the options below as your answer: A. x = 0, –6 B. x = 0, 6 C. x = 6, –6 D. x = 18, –18

OpenStudy (anonymous):

C.) 6, -6

OpenStudy (anonymous):

see yr equation can be put as (x)^2 - (6)^2 then using the formula \[a ^{2}-b ^{2}=(a+b)(a-b)\]

OpenStudy (anonymous):

we get (x)^2 - (6)^2 = 0 ==> (x+6)(x-6) = 0 if x+6 = 0 ==> x = -6 if x -6 = 0 ==> x = 6 so C) 6, -6

jhonyy9 (jhonyy9):

C. is correct